for a wedding, mia bought several dozen of roses and several dozen carnations. the roses cost $15 per dozen, and the carnation cost $8 per dozen. mia bought a total of 17 dozen flowers and paid a total of $192. how much roses did she buy?
Answer by checkley71(8403)

C+R=17 OR C=17-R
15R+8C=192. NOW REPLACE C WITH (17-R) & SOLVE FOR R
15R+8(17-R)=192
15R+136-8R=192
7R=192-136
7R=56
R=56/7
R=8 DOZEN ROSES WERE BOUGHT.
17-8=9 DOZEN CARNATIONS WERE BOUGHT.
PROOF:
15*8+8*9=192
120+72=192
192=192